Dave has $2,000 to invest. He would like to earn $$135.20 per year in interest. How much should he invest at 6% if the rest is to be invested at 7%?

Answer by macston(5194)

.
X=amount at 6%; Y=amount at 7%
X+Y=$2000 Total amount invested is $2000
X=$2000-Y Use this to substitute for X
.
0.06X+0.07Y=$135.20 (Amount of interest at both rates) Substitute for X from above.
0.06($2000-Y)+0.07Y=$135.20
$120-0.06Y+0.07Y=$135.20 Subtract $120 from each side.
0.01Y=$15.20 Multiply each side by 100
Y=$1520.00 He should invest $1520 at 7%.
X=$2000-Y=$2000-$1520=$480 ANSWER: He should invest $480 at 6%.
.
CHECK:
0.06X+0.07Y=$135.20
0.06($480)+0.07($1520)=$135.20
$28.80+$106.40=$135.20
$135.20=$135.20